Find all t in the interval [0, 2π] satisfying (cos t)2 + 10 cos t + 9 = 0.
Answer by lwsshak3(11628)

Find all t in the interval [0, 2π] satisfying (cos t)2 + 10 cos t + 9 = 0
**
Cos^2 t+10cos t+9=0
(cos t+9)(cos t+1)=0
Cos t=-9(reject, (-1 ≤ cos ≤ 1))
or
cos t=-1
t=π